OpenStudy (anonymous):

Widget Wonders produces widgets. they have found that the cost, C(X), of making x wigets is a quadratic function in terms of x. the company also discovers that it costs $30 to produce 2 widgets, $118 to produce 4 widgets, and $766 to produce 10 widgets. find the total cost of producing 6 widgets.

OpenStudy (anonymous):

Use this form to solve your problem \[y=ax ^{2}+bx+c\] Since you have three unknowns, you need three equations to solve the problem. Plug three values given into the general equation. \[30=a2^{2}+b(2) +c\]\[118=a4^{2}=b(4)+c\]\[766=a10^{2}=b(10)+c\] Solve for a, b and c When you find the values for a, b, and c plug them into the general equation with x=6 and solve for y to find your answer.